# Generalization in Reinforcement Learning for Radio Access Networks

### v1

@@ 8.1.1 Distributed Training Architecture
Actors run local replicas of the policy network and interact concurrently with 14 environment instances. [^1] Section 10.2 provides additional computation and deployment details. Similar to Horgan et al. (2018), each actor $i\in\{0,\cdots,39\}$ executes an $\epsilon_{i}$-greedy policy, defined as $\epsilon_{i}=\epsilon^{1+\frac{i}{N-1}\alpha}$, where $\epsilon$ is set to 0.4 and $\alpha$ to 8.5. This exploration policy remains fixed throughout training. On average, each actor generates approximately 132 transitions per second. Incoming transitions are buffered locally before being dispatched asynchronously to replay memory in batches of 500 transitions. Prior to training, a warm-up phase collects 45,000 transitions—without gradient updates—to populate the replay memory with a diverse initial dataset.
@@ 8.1.2 Training Environment Diversification
For hyperparameter tuning, we train RL policies on 40,000 training scenarios generated from 160 random seeds and 250 randomized simulation parameter configurations, as defined in Table 1. Each scenario simulates five seconds (equivalent to 5,000 transmission time intervals), producing approximately 16,000 training samples on average and yielding a total of roughly 650 million samples across the full hyperparameter sweep.

### v2

@@ 8.1.1 Distributed Training Architecture
Actors run local replicas of the policy network and interact concurrently with 14 environment instances (see Section 10.2 that provides additional computation and deployment details). Similar to Horgan et al. (2018), each actor $i\in\{0,\cdots,39\}$ executes an $\epsilon_{i}$-greedy policy, defined as $\epsilon_{i}=\epsilon^{1+\frac{i}{N-1}\alpha}$, where $\epsilon$ is set to 0.4 and $\alpha$ to 8.5. This exploration policy remains fixed throughout training. On average, each actor generates approximately 132 transitions per second. Incoming transitions are buffered locally before being dispatched asynchronously to replay memory in batches of 500 transitions. Prior to training, a warm-up phase collects 45,000 transitions—without gradient updates—to populate the replay memory with a diverse initial dataset.
@@ 8.1.2 Training Environment Diversification
For hyperparameter tuning, we train RL policies on 40,000 training scenarios generated from 160 random seeds and 250 randomized simulation parameter configurations, as defined in Table 1. Each scenario simulates five seconds (equivalent to 5,000 transmission time intervals), producing approximately 16,000 training samples on average and producing a total of approximately 650 million samples across the entire sweep. The resulting hyperparameters are summerized in Section 10.1.